From ear level down the wishing well, a coin is dropped by Liza. The coin falls a distance of 7.00 m before it strikes the water. If the speed of sound is 343 m/s, how long after Liza releases the coin will she hear a splash?


1
Expert's answer
2022-03-08T08:56:19-0500

The total time


"t=t_1+t_2=\sqrt{2h/g}+h/v_s\\
=\rm \sqrt{2*7.00/9.81}+7.00/343=1.22\: s"